If y varies inversely as x, and y = 5 when x = 8, find y when x = 20
Answer by Cromlix(4381)

Hi there,
y varies inversely as x
y = k/x
k is the constant of proportionality.
To find k
y = k/x
Use values: x = 8, y = 5
5 = k/8
Cross multiply
k = 40
Formula:
y = 40/x
Use formula:
y = 40/x
y = 40/20
y = 2.
Hope this helps :-)